# PCLMix: Weakly Supervised Medical Image Segmentation via Pixel-Level Contrastive Learning and Dynamic Mix Augmentation

### v1

@@ 1 Introduction
In this article, we present PCLMix, a robust framework for weakly supervised medical image segmentation designed to address these challenges. PCLMix is structured upon a heterogeneous dual-decoder backbone, incorporating mix augmentation, pixel-level contrastive learning, and dual consistency regularization. For mix augmentation, PCLMix dynamically shuffles image-label pairs within the same batch and mixes them from two different views to generate new samples. These samples are then fed into the network for further training. Subsequently, as illustrated in Figure 1, PCLMix utilizes the outputs of the dual branches to obtain pseudo labels and estimate uncertainty. It strategically selects pixels with high certainty and those within scribble annotations as anchors for pixel-level contrastive learning, thereby reducing noise sampling through uncertainty guidance. Furthermore, we introduce two regularization terms to impose heterogeneous consistency constraints and mixed consistency constraints on model training. To streamline the inference process, both decoders are engaged in training, while during the inference stage, only the main branch (CNN decoder) is utilized for predictions. Our evaluations on the public cardiac dataset ACDC demonstrate the promising performance of PCLMix, achieving comparable or even superior segmentation accuracy when compared to fully supervised approaches. The main contributions of our work are summarized as follows:
@@ 3.1 Overview
In the context of scribble-supervised learning, each sample in the dataset D consists of an image denoted as X and its scribble annotation denoted as $\textbf{Y}=\{(s_{1},l_{1}),(s_{2},l_{2}),\cdots,(s_{K},l_{K})\}$. Here, $K$ is the number of scribble categories, $s_{k}$ refers to the scribble of category $k$, and $l_{k}$ represents the label of category $k$. As shown in Figure 2, the PCLMix architecture is grounded in a one-encoder dual-decoders design, featuring a shared encoder $\mathcal{F}_{e}$ and two independent decoders, i.e., $\mathcal{F}_{c}$ and $\mathcal{F}_{t}$. Specifically, $\mathcal{F}_{c}$ and $\mathcal{F}_{t}$ correspond to the branches constructed by CNN and Transformer, respectively. $\mathcal{F}_{t}$ serves as an auxiliary decoder in the network training and is dropped in the inference phase. This design enables PCLMix to exploit the inherent structural differences between the two decoders without increasing computational complexity.
The proposed PCLMix comprises three key components: dynamic mix augmentation, uncertainty-guided pixel-level contrastive learning, and dual consistency regularization. To effectively propagate sparse supervision signals from local to global in scribble annotations, dynamic mix augmentation is introduced, which expands the sample space by randomly mixing different samples within each batch. However, the distribution of class features in the mixed samples tends to be incomplete and discontinuous, resulting in the potential over-diffusion of supervision signals. To mitigate this issue, uncertainty-guided pixel-level contrastive learning utilizes prediction uncertainty to construct pixel-level contrastive samples, reducing noise sampling and enhancing intra-class consistency and inter-class distinctiveness. Finally, to harness the structural diversity of the heterogeneous dual-branch decoders, dual consistency regularization is employed, which improves the stability of predictions by penalizing inconsistent segmentations between the main decoder and the auxiliary decoder. The overall optimization objective includes the scribble-supervised loss $\ell_{sup}$, contrastive learning loss $\ell_{ctr}$, and consistency regularization loss $\ell_{con}$, which can be formulated as:
@@ 3.3 Uncertainty-Guided Pixel-Level Contrastive Learning
Inspired by [29], we utilize contrastive learning in a low-resolution feature space to improve computational efficiency and capture more extensive semantic information. To efficiently manage computational overhead, we establish a memory queue to store acquired samples, holding prototype vectors alongside their corresponding pixel categories. The widely used InfoNCE [4] loss function for computing contrastive loss is employed. In each iteration, a random selection of $N$ anchor points is made, and the contrastive loss is calculated for each anchor. The overall contrastive loss is then determined by averaging the losses from all selected anchors. The calculation process is illustrated as follows:
@@ 3.4 Regularization of Supervision via Dual Consistency
Here, $\textbf{p}^{12}$ and $\textbf{p}^{21}$ are respectively the mixed predictions of the new images $\textbf{x}^{12}$ and $\textbf{x}^{21}$. Following previous work [35, 8], we define the mixed consistency loss as the negative cosine similarity between two segmentation results:
@@ 4.1 Experimental Setup
Table 1: Comparison of our method with state-of-the-art scribble-supervised approaches on the ACDC dataset. All evaluations are conducted through 5-fold cross-validation without utilizing ensemble strategies. The average results (standard deviation) for other methods are sourced from [37]. Bold indicates the best performance, while underline signifies the second-best. $\uparrow$ denotes lower values are better, $\downarrow$ indicates higher values are preferred.
@@ 4.2 Quantitative Comparison
Table LABEL:quantComp presents the results of our quantitative comparative experiments on the ACDC dataset. Clearly, compared to existing methods, PCLMix demonstrates more promising outcomes. In terms of improvement magnitude, PCLMix outperforms the baseline significantly on the ACDC dataset, increasing the Dice score by 20.1 and reducing the HD95 by 169.2. Furthermore, when compared with state-of-the-art methods (SOTAs), PCLMix achieves the best Dice performance and the best HD95 performance among all scribble-supervised approaches. This indicates that, in overall segmentation performance, PCLMix exhibits slight enhancement compared to SOTAs while achieving substantial improvement in boundary segmentation. It’s worth noting that, compared to the upper limit, our method exhibits a slight decrease in Dice (only 1.1) but surprisingly good results in HD95, with a relative reduction of 41.4%. This shows that our approach successfully overcomes the inherent constraints associated with sparse supervision, narrowing or even eliminating the gap with fully supervised methods.
@@ 4.3 Ablation Study
In version #1, with only supervised loss, baseline performance is observed. Version #7, incorporating all loss components, achieves peak performance with a Dice score of 88.9 and HD95 of 4.1. Comparison between versions #1, #3, and #4 highlights the significant contribution of dual consistency regularization to improving model performance. Additionally, analyzing versions #1, #2, #4, and #7 reveals a limited impact of contrastive learning during poor model performance, but it proves beneficial under optimal conditions. We speculate that this is because, even with the guidance of uncertainty maps, contrastive learning still introduces noise pixels in situations of poor performance. The performance drop in version #6, compared to #7, emphasizes the Transformer’s crucial role as an auxiliary decoder.

### v2

@@ 1 Introduction
In this article, we present PCLMix, a robust framework for weakly supervised medical image segmentation designed to address these challenges. PCLMix is structured upon a heterogeneous dual-decoder backbone, incorporating mix augmentation, pixel-level contrastive learning, and dual consistency regularization. For mix augmentation, PCLMix dynamically shuffles image-label pairs within the same batch and mixes them from two different views to generate new samples. These samples are then fed into the network for further training. Subsequently, as illustrated in Figure 1 PCLMix utilizes the outputs of the dual branches to obtain pseudo labels and estimate uncertainty. It strategically selects pixels with high certainty and those within scribble annotations as anchors for pixel-level contrastive learning, thereby reducing noise sampling through uncertainty guidance. Furthermore, we introduce two regularization terms to impose heterogeneous consistency constraints and mixed consistency constraints on model training. To streamline the inference process, both decoders are engaged in training, while during the inference stage, only the main branch (CNN decoder) is utilized for predictions. Our evaluations on the public cardiac dataset ACDC demonstrate the promising performance of PCLMix, achieving comparable or even superior segmentation accuracy when compared to fully supervised approaches. The main contributions of our work are summarized as follows:
@@ 3.1 Overview
In scribble-supervised learning, each sample in the dataset D includes an image denoted as X and its corresponding scribble annotation denoted as $\textbf{Y}={(s_{1},l_{1}),(s_{2},l_{2}),\cdots,(s_{K},l_{K})}$, where $K$ represents the number of scribble categories, $s_{k}$ refers to the scribble of category $k$, and $l_{k}$ represents the label of category $k$. Figure 2 illustrates the PCLMix architecture, which is based on a one-encoder dual-decoders design featuring a shared encoder $\mathcal{F}{e}$ and two independent decoders, denoted as $\mathcal{F}{c}$ (CNN-based) and $\mathcal{F}{t}$ (Transformer-based). While $\mathcal{F}{t}$ serves as an auxiliary decoder during network training, it is not used during the inference phase. This design allows PCLMix to leverage the structural differences between the decoders without increasing computational complexity.
PCLMix consists of three main components: dynamic mix augmentation, uncertainty-guided pixel-level contrastive learning, and dual consistency regularization. Dynamic mix augmentation expands the sample space by randomly mixing different samples within each batch, aiding in the propagation of sparse supervision signals from local to global regions in scribble annotations. However, mixed samples may have incomplete and discontinuous class feature distributions, potentially causing over-diffusion of supervision signals. To address this, uncertainty-guided pixel-level contrastive learning uses prediction uncertainty to construct contrastive samples at the pixel level, reducing noise sampling and improving intra-class consistency and inter-class distinctiveness. Finally, dual consistency regularization leverages the structural diversity of the heterogeneous dual-branch decoders, enhancing prediction stability by penalizing inconsistent segmentations between the main and auxiliary decoders. The overall optimization objective encompasses the scribble-supervised loss $\ell_{sup}$, contrastive learning loss $\ell_{ctr}$, and consistency regularization loss $\ell_{con}$, formulated as follows:
@@ 3.3 Uncertainty-Guided Pixel-Level Contrastive Learning
Inspired by [29], we utilize contrastive learning in a low-resolution feature space to improve computational efficiency and capture more extensive semantic information. To efficiently manage computational overhead, we establish a memory queue to store acquired samples, holding prototype vectors alongside their corresponding pixel categories. The widely used InfoNCE loss function for computing contrastive loss is employed. In each iteration, a random selection of $N$ anchor points is made, and the contrastive loss is calculated for each anchor. The overall contrastive loss is then determined by averaging the losses from all selected anchors. The calculation process is illustrated as follows:
@@ 3.4 Regularization of Supervision via Dual Consistency
Here, $\textbf{p}^{12}$ and $\textbf{p}^{21}$ are respectively the mixed predictions of the new images $\textbf{x}^{12}$ and $\textbf{x}^{21}$. Following previous work [35, Chen_He_2021, 8], we define the mixed consistency loss as the negative cosine similarity between two segmentation results:
@@ 4.1 Experimental Setup
Table 1: Comparison of our method with state-of-the-art scribble-supervised approaches on the ACDC dataset. All evaluations are conducted through a 5-fold cross-validation. The average results (standard deviation) for other methods are sourced from [28]. Bold indicates the best performance, while underline signifies the second-best. $\uparrow$ denotes lower values are better, $\downarrow$ indicates higher values are preferred, and marks methods employing an ensemble strategy.
@@ 4.2 Quantitative Comparison
Table LABEL:quantComp presents the results of our quantitative comparative experiments on the ACDC dataset. Clearly, compared to existing methods, PCLMix demonstrates more prom-ising outcomes. In terms of improvement magnitude, PCLMix outperforms the baseline significantly on the ACDC dataset, increasing the Dice score by 20.1 and reducing the HD95 by 169.2. Furthermore, when compared with state-of-the-art methods (SOTAs), PCLMix achieves the second-best Dice performance and the best HD95 performance among all scribble-supervised approaches. This indicates that, in overall segmentation performance, PCLMix is on par with SOTAs while achieving substantial improvement in boundary segmentation. It’s worth noting that, compared to the upper limit, our method exhibits a slight decrease in Dice (only 1.1) but surprisingly good results in HD95, with a relative reduction of 41.4%. This shows that our approach successfully overcomes the inherent con-straints associated with sparse supervision, narrowing or even eliminating the gap with fully supervised methods.
@@ 4.3 Ablation Study
In version #1, with only supervised loss, baseline performance is observed. Version #7, incorporating all loss components, achieves peak performance with a Dice score of 88.7 and HD95 of 4.1. Comparison between versions #1, #3, and #4 highlights the significant contribution of dual consistency regularization to improving model performance. Additionally, analyzing versions #1, #2, #4, and #7 reveals a limited impact of contrastive learning during poor model performance, but it proves beneficial under optimal conditions. We speculate that this is because, even with the guidance of uncertainty maps, contrastive learning still introduces noise pixels in situations of poor performance. The performance drop in version #6, compared to #7, emphasizes the Transformer’s crucial role as an auxiliary decoder.
@@ 4.5 Visualization Analysis
The experiments demonstrate exciting improvements of PCLMix in HD95, suggesting its superior segmentation performance, particularly in boundary regions, compared to the SOTAs, which can be further supported by Figure 4. Furthermore, although PCLMix lags behind TriMix in Dice scores, it significantly reduces computational complexity and improves inference efficiency by discarding the auxiliary decoder branch during inference.
However, we also observed that introducing the proposed contrastive learning strategy may lead to further performance degradation, especially when the network’s performance is poor. We attribute this to the inevitable introduction of more noisy pixels during anchor sampling, even with guidance from uncertainty maps. Lastly, due to the presence of the auxiliary decoder, PCLMix exhibits higher computational complexity during the training phase, limiting the exploration of ensemble methods due to resource constraints. We hope to address this limitation in future research endeavors.